# Bright traveling breathers in media with long-range, nonconvex dispersion

### v2

@@ I INTRODUCTION
Coherently propagating disturbances such as solitary waves and envelope solitary waves, ubiquitous in nonlinear dispersive media, are formed due to a balance between nonlinearity and dispersion. Envelope solitary waves are generically described by solutions of the cubic nonlinear Schrödinger (NLS) equation subject to two disparate spatial scales: the fast spatial scale corresponding to the wavelength of periodic carrier wave oscillations and the slow amplitude/phase modulation scale. For an attractive, focusing nonlinear medium, the NLS equation admits bright soliton solutions corresponding to a localized, sech envelope modulation of a rapidly varying carrier wave. When the carrier phase speed and envelope group speed differ, NLS bright solitons approximate unsteady nonlinear wave packets. Such nonlinear wavepackets are also referred to as breathers due to the pulsation or breathing of their internal oscillations. But the mere existence of approximate NLS bright soliton solutions does not guarantee the existence of breather solutions to a nonlinear dispersive evolution equation.
The canonical evolution equations admitting breather solutions are the sine-gordon (SG) Ablowitz and Segur (1981) and modified Korteweg-de Vries (mKdV) Clarke et al. (2000) equations. These equations are completely integrable by the inverse scattering transform where soliton and breather solutions correspond to discrete eigenvalues of a corresponding linear spectral problem Ablowitz and Segur (1981). A breather solution of SG or mKdV corresponds to a pair or quartet of eigenvalues, respectively, which can be interpreted as a bound state of two solitons that decay to zero Clarke et al. (2000). Breather solutions to the focusing NLS equation have also been identified as important for modeling rogue wave phenomena Onorato et al. (2013) and exhibit a nonzero, plane wave background.
@@
Such breathers and traveling breathers are also prevalent, for example, in fluid dynamics Lamb et al. (2007); Grimshaw and Helfrich (2008); Whitfield and Johnson (2015); Maiden and Hoefer (2016); Nakayama and Lamb (2020), nonlinear and matter-wave optics Leblond and Mihalache (2013); Hasegawa and Kodama (1995); Kivshar and Luther-Davies (1998); Mandelik et al. (2003); Schäfer and Wayne (2004); Sakovich and Sakovich (2005); Di Carli et al. (2019); Luo et al. (2020), and magnetic materials Kalinikos et al. (1990); Ustinov et al. (2008); Drozdovskii et al. (2010). However, due to their inherently unsteady character, breather and traveling breather solutions are challenging to obtain when exact solution methods fail, as is typical in many physical model equations which are not integrable. As noted earlier, the exact breather waveforms for integrable systems are constructed through a nonlinear “superposition” principle based on an application of the Darboux transformation. On the other hand, a common approach to circumvent the reduced analytical tractability associated with nonintegrable systems involves long-time numerical evolution of suitably chosen initial conditions that appear to lead to breather solutions Helfrich (2007); Lamb et al. (2007); Grimshaw and Helfrich (2008); Dyachenko and Zakharov (2008); Whitfield and Johnson (2015); Maiden and Hoefer (2016); Nakayama and Lamb (2020). But, due to the existence of small amplitude radiation accompanying the time evolution, it is difficult to discern whether breather solutions actually exist and, if so, what their properties are. The existence of radiation in these and other computational studies Kudryavtsev (1975); Ablowitz et al. (1979); Campbell et al. (1983) suggests that a more likely scenario for non-integrable systems is that breathers are accompanied by an oscillatory background Segur and Kruskal (1987); Boyd (1990); Soffer and Weinstein (1999), recent numerical evidence of localized breathers in a nonintegrable equation notwithstanding Kalisch et al. (2022). In other words, just as solitary wave solutions in non-integrable equations generalize the soliton solutions of integrable equations, traveling breather solutions of non-integrable equations are the natural generalization of breather solutions of integrable equations.
@@
In the present study, we numerically investigate the existence of bright (elevation) traveling breathers to the Benjamin-Bona-Mahony (BBM) and conduit equations, both nonintegrable, non-evolutionary equations Mikhailov et al. (2007), by solving a space-time boundary value problem (BVP) in the co-moving reference frame where the envelope speed is zero. Periodic in time together with periodic in space boundary conditions are enforced. Multiple one-dimensional families of traveling breather solutions are obtained by numerically continuing the BVP solutions from the weakly nonlinear Schrödinger bright soliton approximation with a given carrier wavenumber. The unit-mean carrier frequency, phase shift, and amplitude are implicitly determined by fixing the breather velocity, the carrier frequency in the co-moving frame, and the spatial domain size. Solution branches are obtained by performing continuation in the traveling breather velocity, which is negative for all solution branches computed.
@@
Solutions along a given continuation branch are found to strongly depend upon the initial carrier wavenumber. When the wavenumber is sufficiently far from the inflection point of the linear dispersion relation, the carrier background amplitude grows with decreasing velocity while the traveling breather width narrows relative to the carrier wavelength. For initial carrier wavenumbers close to the inflection point, the traveling breather envelope width remains large relative to the carrier wavelength. The traveling breather solutions obtained here are found to be dynamically stable under long-time numerical evolution subject to small amplitude initial noise. Since nonlinear short-pulse optics Costanzino et al. (2009); Leblond and Mihalache (2013) and internal oceanic waves influenced by the earth’s rotation Helfrich (2007); Whitfield and Johnson (2015) exhibit similar nonconvex, rational dispersion, the traveling breathers obtained in our study may have implications for these and other applications.
@@ II Model equation properties
is a long-wavelength model of weakly nonlinear waves. Equation (1) is not integrable, possessing exactly 2 other linearly independent conservation laws Olver (1979). Besides the usual space and time translational invariance properties, the BBM possesses the scaling symmetry
@@
where $u_{0}$ is a nonzero, real constant. The BBM equation’s linear dispersion relation for trigonometric traveling waves on the constant background $u_{0}$ is bounded
${\omega_{0}({k},u_{0})=\frac{u_{0}{k}}{1+{k}^{2}},}$ (3)
@@
The BBM equation (1) admits a three-parameter family of periodic traveling wave solutions in the form of cnoidal waves
$\begin{split}u(x,t)&=\tilde{\beta}+(\gamma-\tilde{\beta})\mathrm{cn}^{2}\left(z,m\right),\\ z&=\left(\frac{\gamma-\alpha}{12s}\right)^{1/2}(x-st),\quad m=\frac{\gamma-\tilde{\beta}}{\gamma-\alpha},\end{split}$ (5)
where $\alpha<\tilde{\beta}<\gamma$, $s=\frac{1}{3}(\alpha+\tilde{\beta}+\gamma)$ is the phase velocity and $\mathrm{cn}(z,m)$ is the Jacobi elliptic cosine function. The cnoidal wave’s amplitude $a$ and wavenumber $k$ are
$a=\gamma-\tilde{\beta},\quad k=\frac{2\pi}{L},\quad L=4K(m)\sqrt{\frac{3s}{\gamma-\alpha}},$ (6)
@@
$\overline{u}=\alpha+(\gamma-\alpha)\frac{E(m)}{K(m)},$ (7)
where $K(m)$ and $E(m)$ are the complete elliptic integrals of the first and second kinds, respectively. The cnoidal wave (5) limits to a solitary wave when $\tilde{\beta}\to\alpha$ and a constant when $\tilde{\beta}\to\gamma$.
By use of the scaling symmetry (2), we impose the unit mean constraint $\overline{u}=1$ on the cnoidal wave solutions without loss of generality, and therefore constrain $\alpha$, $\tilde{\beta}$, and $\gamma$ via $\overline{u}=1$ in eq. (7). We parameterize the set of unit mean periodic traveling wave solutions to the BBM equation in terms of two parameters such as $(a,k)$. Then, its frequency is determined $\omega=ks$.
@@
modeling large amplitude, long waves along the circular, free interface between two viscous fluids with high viscosity contrast and small Reynolds number Lowman and Hoefer (2013). Its linear dispersion relation on the constant background $A_{0}>0$ is similar to BBM’s (3)
@@
The conduit equation (8) also admits a three-parameter family of periodic traveling wave solutions Maiden and Hoefer (2016); Johnson and Perkins (2020) but an analytical expression for it is not known. Utilizing the scaling symmetry (10), we impose the unit mean constraint so that the cnoidal-like periodic traveling wave solutions are parameterized by, e.g., their wavenumber $k$ and amplitude $a$.
@@ III Computational methodology
We begin by introducing the common approach to computing approximate bright breather solutions by assuming weak nonlinearity and scale separation. The NLS equation models the slowly varying envelope, $B(\zeta,\tilde{\tau})$, of nearly monochromatic nonlinear wavepackets Ablowitz (2011); Whitham (2011). In this regard, the focusing NLS reduction obtained by employing a standard multiple scales calculation and rescaling takes the form
@@
where $\partial_{\tilde{k}}\omega_{0}$ is the group velocity, $\partial_{\tilde{k}\tilde{k}}\omega_{0}$ represents the dispersion curvature and $\epsilon$ is an amplitude scale. The benefit of the NLS approximation is that a simple, ordinary differential equation (ODE) can be sought for describing the spatial variation of the envelope $b(\zeta)$, where $B=b(\zeta)\exp(i\mu\tilde{\tau})$ and $\mu$ is an amplitude dependent frequency shift. The well-known sech-solution of this ODE $b\equiv{\rm sech}(\zeta),\mu=1/2$ will be used to seed the continuation algorithm that we describe shortly.
In contrast, the direct computation of traveling breathers requires solving a partial differential equation (PDE). We now describe the strategy we adopt to compute BBM and conduit bright traveling breathers as solutions to a space-time boundary value problem.
@@ III.1 Space-time boundary value problem
where $\mathbb{L}$ is a linear, constant coefficient skew-adjoint differential operator while $\mathbb{N}[u,u_{t}]$ is in general a nonlinear operator acting on $u$ and $u_{t}$. Entering the co-moving frame with velocity $c$ $(\chi=x-ct,\tau=t)$, we recast eq. (13) as
@@
where, $u(x,t)\equiv 1+\beta(\chi,\tau)$ and the linear operator $\tilde{\mathbb{L}}$ inherits the skew-symmetry. Using the scaling symmetry (2) or (10), we set the background mean to unity, without loss of generality. In this reference frame, the solution is assumed to be time-periodic with period $T$ and to rapidly decay to a periodic background in space. Then, $\beta$ has zero mean in the far field in $\chi$ due to the unit mean normalization of $u$. We truncate the domain $(\chi,\tau)\in[-L,L]\times[0,T]$ and take a finite Fourier product basis for the solution field
$\beta(\chi,\tau;c,T)$ $\approx\sum_{m=-N}^{N}\alpha_{m}(\chi)e^{im\frac{2\pi\tau}{T}}$ (15)
$\alpha_{m}(\chi)$ $\approx\sum_{s=-M}^{M}\hat{\alpha}_{ms}e^{is\frac{2\pi\chi}{L}}.$
The spatial domain is chosen to be sufficiently large so that boundary effects are negligible, leading to a fully periodic product basis. The action of the two-dimensional forward and inverse, discrete Fourier transform is denoted by $\mathcal{F}_{2D}(\cdot)$ and $\mathcal{F}_{2D}^{-1}(\cdot)$, respectively.
We now describe the iterative procedure to recover numerical solutions from their weakly nonlinear approximations governed by the NLS equation, which in the co-traveling frame, and in terms of fast space and time variables is :
@@
where $\epsilon=\frac{\tilde{a}\sqrt{\gamma}}{4}$ while the parameters $\gamma(\tilde{k})$ and $\omega_{0}$ depend on the particular dispersive evolution equation at hand. Having fixed the time-period $T$ to the NLS prediction, the solution field $\beta$ and the envelope velocity $c$ need to be determined. As we will show, the family of unit-mean solutions is three dimensional. While the ansatz (15) for $\beta$ can be used in Eq. (14), to determine the velocity $c$, we require an additional condition. Multiplying Eq. (14) by $\beta_{\tau}$, and isolating the terms containing the velocity $c$ results in $\mathcal{H}(\beta)=c\mathcal{G}(\beta)$. Integrating this expression over the entire spatio-temporal domain, we obtain the following self-consistent integral condition for $c$
@@
We implement the Newton-conjugate gradient algorithm on eq. (14) subject to the ansatz (15). To recover $c$ and $\beta$ simultaneously, we update the velocity iteratively at every outer Newton iteration using (17) (see Yang (2015) for a similar treatment). Upon insertion of (15) into (14), Newton operator iterations are then applied to Eq. (14). The linearization is symmetrized and inner, preconditioned-conjugate gradient iterations are used to solve this positive, semi-definite self-adjoint system, which is expected to converge Yang (2009). The complete velocity-solution profile update algorithm is summarized as
${\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{1n}\Delta\beta$ $=-{\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{0}\beta_{n};n=1,2,\cdots$ (18)
@@
where ${\textbf{P}}_{1n}$ is the linearization operator at the $n^{th}$ Newton iteration, with its adjoint operator represented by ${\textbf{P}}_{1n}^{\dagger}$, and ${\bf P}_{0}\beta_{n}$ is the residual of Eq. (14) at the $n^{th}$ iteration.
We seed the iterations with a sufficiently close initial guess ($\beta_{1},c_{1}$). For the BBM and conduit equations, the bright solitary wave solutions of the corresponding NLS reductions give sufficiently close approximations. Iterations are terminated when the residual $\underset{\chi,\tau}{\rm max}\;|{\bf P}_{0}\beta|$ is less than $10^{-10}$. To reduce the condition number of the linear operator ${\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{1n}$, it is necessary to introduce an acceleration operator. We follow the guidelines outlined in Yang (2009, 2015). For Eq. (14), an appropriate acceleration operator at the $n^{th}$ Newton iteration is chosen by examining the constant-coefficient part of the symmetrized operator ${\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{1n}$, given by $-(\mathbb{B}_{n})^{2}=-(\partial_{\tau}-c_{n}\partial_{\chi}-\tilde{\mathbb{L}})^{2}$. Notably, this operator is positive semi-definite. An appropriate acceleration operator is thus given by the positive definite operator ${\bf M}_{n}=r-(\mathbb{B}_{n})^{2}$, where $r>0$ is a positive number whose optimal choice is arrived at via numerical experiments.

### v3

@@ I INTRODUCTION
Coherently propagating disturbances such as solitary waves and envelope solitary waves, ubiquitous in nonlinear dispersive media, are formed due to a balance between nonlinearity and dispersion. Envelope solitary waves are generically described by solutions of the cubic nonlinear Schrödinger (NLS) equation subject to two disparate spatial scales: the fast spatial scale corresponding to the wavelength of periodic carrier wave oscillations and the slow-amplitude-phase-modulation scale. For an attractive, focusing nonlinear medium, the NLS equation admits bright soliton solutions corresponding to a localized, sech envelope modulation of a rapidly varying carrier wave. When the carrier phase speed and envelope group speed differ, NLS bright solitons approximate unsteady nonlinear wave packets. Such nonlinear wavepackets are also referred to as breathers due to the pulsation or breathing of their internal oscillations. However, the mere existence of approximate NLS bright soliton solutions does not guarantee the existence of breather solutions to a nonlinear dispersive evolution equation.
The canonical evolution equations admitting breather solutions are the sine-Gordon (SG) Ablowitz and Segur (1981) and modified Korteweg-de Vries (mKdV) Clarke et al. (2000) equations. These equations are completely integrable by the inverse scattering transform where soliton and breather solutions correspond to discrete eigenvalues of a corresponding linear spectral problem Ablowitz and Segur (1981). A breather solution of SG or mKdV equations corresponds to a pair or quartet of eigenvalues, respectively, which can be interpreted as a bound state of two solitons that decay to zero Clarke et al. (2000). Breather solutions to the focusing NLS equation have also been identified as important for modeling rogue wave phenomena Onorato et al. (2013) and exhibit a nonzero, plane wave background.
@@
Such breathers and traveling breathers are also prevalent, for example, in fluid dynamics Lamb et al. (2007); Grimshaw and Helfrich (2008); Whitfield and Johnson (2015); Maiden and Hoefer (2016); Nakayama and Lamb (2020), nonlinear and matter-wave optics Leblond and Mihalache (2013); Hasegawa and Kodama (1995); Kivshar and Luther-Davies (1998); Mandelik et al. (2003); Schäfer and Wayne (2004); Sakovich and Sakovich (2005); Di Carli et al. (2019); Luo et al. (2020), and magnetic materials Kalinikos et al. (1990); Ustinov et al. (2008); Drozdovskii et al. (2010). However, due to their inherently unsteady character, breather and traveling breather solutions are challenging to obtain. As noted earlier, the exact breather waveforms for integrable systems are constructed through a nonlinear “superposition” principle based on an application of the Darboux transformation. On the other hand, a common approach to circumvent the reduced analytical tractability associated with nonintegrable systems involves long-time numerical evolution of suitably chosen initial conditions that appear to lead to breather solutions Helfrich (2007); Lamb et al. (2007); Grimshaw and Helfrich (2008); Dyachenko and Zakharov (2008); Whitfield and Johnson (2015); Maiden and Hoefer (2016); Nakayama and Lamb (2020). However, due to the existence of small-amplitude radiation accompanying the time evolution, it is difficult to discern whether breather solutions actually exist and, if so, what their properties are. The existence of radiation in these and other computational studies Kudryavtsev (1975); Ablowitz et al. (1979); Campbell et al. (1983) suggests that a more likely scenario for nonintegrable systems is that breathers are accompanied by an oscillatory background Segur and Kruskal (1987); Boyd (1990); Soffer and Weinstein (1999), recent numerical evidence of localized breathers in a nonintegrable equation notwithstanding Kalisch et al. (2022). In other words, just as solitary-wave solutions in nonintegrable equations generalize the soliton solutions of integrable equations, traveling breather solutions of nonintegrable equations are the natural generalization of breather solutions of integrable equations.
@@
In the present study, we numerically investigate the existence of bright (elevation) traveling breathers to the Benjamin-Bona-Mahony (BBM) and conduit equations, both nonintegrable, nonevolutionary equations Mikhailov et al. (2007), by solving a space-time boundary-value problem (BVP) in the comoving reference frame where the envelope speed is zero. The boundary conditions are periodicity in time and space. Multiple one-dimensional families of traveling breather solutions are obtained by numerically continuing the BVP solutions from the weakly nonlinear Schrödinger bright soliton approximation with a given carrier wave number and amplitude. The unit-mean carrier frequency, phase shift, and amplitude are implicitly determined by fixing the breather velocity, the carrier frequency in the comoving frame, and the spatial domain size. Solution branches are obtained by performing continuation in the traveling breather velocity, which is negative for all solution branches computed.
@@
Solutions along a given continuation branch are found to strongly depend upon the initial carrier wave number. When the wave number is sufficiently far from the inflection point of the linear dispersion relation, the carrier background amplitude grows with decreasing velocity while the traveling breather width narrows relative to the carrier wavelength. For initial carrier wave numbers close to the inflection point, the traveling breather envelope width remains large relative to the carrier wavelength. The traveling breather solutions obtained here are found to be dynamically stable under long-time numerical evolution subject to small-amplitude initial noise. Since nonlinear short-pulse optics Costanzino et al. (2009); Leblond and Mihalache (2013) and internal oceanic waves influenced by the earth’s rotation Helfrich (2007); Whitfield and Johnson (2015) exhibit similar nonconvex rational dispersion, the traveling breathers obtained in our study may have implications for these and other applications.
@@ II Model equation properties
is a long-wavelength model of weakly nonlinear waves. Equation (LABEL:BBM-equation-1) is not integrable, possessing exactly two other linearly independent conservation laws Olver (1979). Besides the usual space and time translational-invariance properties, the BBM possesses the scaling symmetry
@@
where $u_{0}$ is a nonzero real constant. The BBM equation’s linear dispersion relation for trigonometric traveling waves on the constant background $u_{0}$ is bounded
${\omega_{0}({k},u_{0})=\frac{u_{0}{k}}{1+{k}^{2}}}$ (3)
@@
The BBM equation (LABEL:BBM-equation-1) admits a three-parameter family of periodic traveling-wave solutions in the form of cnoidal waves
$\begin{split}u(x,t)&=\tilde{\beta}+(\tilde{\gamma}-\tilde{\beta})\mathrm{cn}^{2}\left(z,m\right),\\ z&=\left(\frac{\tilde{\gamma}-\tilde{\alpha}}{12\tilde{s}}\right)^{1/2}(x-\tilde{s}t),\quad m=\frac{\tilde{\gamma}-\tilde{\beta}}{\tilde{\gamma}-\tilde{\alpha}},\end{split}$ (5)
where $\tilde{\alpha}<\tilde{\beta}<\tilde{\gamma}$, $\tilde{s}=\frac{1}{3}(\tilde{\alpha}+\tilde{\beta}+\tilde{\gamma})$ is the phase velocity and $\mathrm{cn}(z,m)$ is the Jacobi elliptic cosine function. The cnoidal wave’s amplitude $a$ and wavenumber $k$ are
$a=\tilde{\gamma}-\tilde{\beta},\quad k=\frac{2\pi}{L},\quad L=4K(m)\sqrt{\frac{3\tilde{s}}{\tilde{\gamma}-\tilde{\alpha}}},$ (6)
@@
$\overline{u}=\tilde{\alpha}+(\tilde{\gamma}-\tilde{\alpha})\frac{E(m)}{K(m)},$ (7)
where $K(m)$ and $E(m)$ are the complete elliptic integrals of the first and second kinds, respectively. The cnoidal wave (5) limits to a solitary wave when $\tilde{\beta}\to\tilde{\alpha}$ and a constant when $\tilde{\beta}\to\tilde{\gamma}$.
By use of the scaling symmetry (2), we impose the unit-mean constraint $\overline{u}=1$ on the cnoidal wave solutions without loss of generality and therefore constrain $\tilde{\alpha}$, $\tilde{\beta}$, and $\tilde{\gamma}$ via $\overline{u}=1$ in eq. (7). We parametrize the set of unit mean periodic traveling-wave solutions to the BBM equation in terms of two parameters such as $(a,k)$. Then its frequency is determined to be $\omega=k\tilde{s}$.
@@
modeling large amplitude, long waves along the circular, free interface between two viscous fluids with high viscosity contrast and small Reynolds number Lowman and Hoefer (2013). Its linear dispersion relation on the constant background $A_{0}>0$ is similar to the BBM equation (3)
@@
The conduit equation (LABEL:Conduit-equation-1) also admits a three-parameter family of periodic traveling-wave solutions Maiden and Hoefer (2016); Johnson and Perkins (2020) but an analytical expression for it is not known. Utilizing the scaling symmetry (10), we impose the unit-mean constraint so that the cnoidal-like periodic traveling-wave solutions are parametrized by, e.g., their wave number $k$ and amplitude $a$.
@@ III Computational methodology
We begin by introducing the common approach to computing approximate bright breather solutions by assuming weak nonlinearity and scale separation. The NLS equation models the slowly varying envelope $B(\zeta,\tilde{\tau})$ of nearly monochromatic nonlinear wavepackets Ablowitz (2011); Whitham (2011). In this regard, the focusing NLS reduction obtained by employing a standard multiple-scales calculation and rescaling takes the form
@@
where $\partial_{\tilde{k}}\omega_{0}$ is the group velocity, $\partial_{\tilde{k}\tilde{k}}\omega_{0}$ represents the dispersion curvature, and $\epsilon$ is an amplitude scale. The benefit of the NLS approximation is that a simple ordinary differential equation (ODE) can be sought for describing the spatial variation of the envelope $b(\zeta)$, where $B=b(\zeta)\exp(i\mu\tilde{\tau})$ and $\mu$ is an amplitude-dependent frequency shift. The well-known sech solution of this ODE, $b\equiv{\rm sech}(\zeta)$ and $\mu=\frac{1}{2}$ will be used to seed the continuation algorithm that we describe shortly.
In contrast, the direct computation of traveling breathers requires solving a partial differential equation. We now describe the strategy we adopt to compute BBM and conduit bright traveling breathers as solutions to a space-time boundary-value problem.
@@ III.1 Space-time boundary value problem
where $\mathbb{L}$ is a linear, constant coefficient skew-adjoint differential operator while $\mathbb{N}[u,u_{t}]$ is in general a nonlinear operator acting on $u$ and $u_{t}$. Entering the comoving frame with velocity $c$ $(\chi=x-ct,\tau=t)$, we recast eq. (13) as
@@
where $u(x,t)\equiv 1+\beta(\chi,\tau)$ and the linear operator $\tilde{\mathbb{L}}$ inherits the skew symmetry. Using the scaling symmetry (2) or (10), we set the background mean to unity, without loss of generality. In this reference frame, the solution is assumed to be time periodic with period $T$ and to rapidly decay to a periodic background in space. Then $\beta$ has zero mean in the far field in $\chi$ due to the unit-mean normalization of $u$. We truncate the domain $(\chi,\tau)\in[-L,L]\times[0,T]$ and take a finite Fourier product basis for the solution field
$\beta(\chi,\tau;c,T)$ $\approx\sum_{m=-N}^{N}\alpha_{m}(\chi)e^{im({2\pi\tau}/{T})},$ (15)
$\alpha_{m}(\chi)$ $\approx\sum_{s=-M}^{M}\hat{\alpha}_{ms}e^{is({2\pi\chi}/{L})}.$
The spatial domain is chosen to be sufficiently large so that boundary effects are negligible, leading to a fully periodic product basis. The actions of the two-dimensional forward and inverse discrete Fourier transforms are denoted by $\mathcal{F}_{2{\rm D}}(\cdot)$ and $\mathcal{F}_{2{\rm D}}^{-1}(\cdot)$, respectively.
We now describe the iterative procedure to recover numerical solutions from their weakly nonlinear approximations governed by the NLS equation, which in the co-traveling frame, and in terms of fast space and time variables is
@@
where $\epsilon=\frac{\tilde{a}\sqrt{\gamma}}{4}$ while the parameters $\gamma(\tilde{k})$ and $\omega_{0}$ depend on the particular dispersive evolution equation at hand. Having fixed the time period $T$ to the NLS prediction, the solution field $\beta$ and the envelope velocity $c$ need to be determined. As we will show, the family of unit-mean solutions is three dimensional. While the ansatz (15) for $\beta$ can be used in Eq. (14), to determine the velocity $c$, we require an additional condition. Multiplying Eq. (14) by $\beta_{\tau}$ and isolating the terms containing the velocity $c$ results in $\mathcal{H}(\beta)=c\mathcal{G}(\beta)$. Integrating this expression over the entire spatiotemporal domain, we obtain the self-consistent integral condition for $c$,
@@
We implement the Newton-conjugate gradient algorithm on eq. (14) subject to the ansatz (15). To recover $c$ and $\beta$ simultaneously, we update the velocity iteratively at every outer Newton iteration using (17) (see Yang (2015) for a similar treatment). Upon insertion of (15) into (14), Newton operator iterations are then applied to Eq. (14). The linearization is symmetrized and inner preconditioned-conjugate gradient iterations are used to solve this positive-semidefinite self-adjoint system, which is expected to converge Yang (2009). The complete velocity-solution profile update algorithm is summarized as
${\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{1n}\Delta\beta$ $=-{\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{0}\beta_{n},\;\;\;n=1,2,_{\cdots}$ (18)
@@
where ${\textbf{P}}_{1n}$ is the linearization operator at the $n$th Newton iteration, with its adjoint operator represented by ${\textbf{P}}_{1n}^{\dagger}$, and ${\bf P}_{0}\beta_{n}$ is the residual of Eq. (14) at the $n$th iteration.
We seed the iterations with a sufficiently close initial guess ($\beta_{1}\;{\rm and}\;c_{1}$). For the BBM and conduit equations, the bright solitary-wave solutions of the corresponding NLS reductions give sufficiently close approximations. Iterations are terminated when the residual $\underset{\chi,\tau}{\rm max}\;|{\bf P}_{0}\beta|$ is less than $10^{-10}$. To reduce the condition number of the linear operator ${\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{1n}$, it is necessary to introduce an acceleration operator. We follow the guidelines outlined in Yang (2009, 2015). For Eq. (14), an appropriate acceleration operator at the $n$th Newton iteration is chosen by examining the constant-coefficient part of the symmetrized operator ${\textbf{P}}_{1n}^{\dagger}{\textbf{P}}_{1n}$, given by $-(\mathbb{B}_{n})^{2}=-(\partial_{\tau}-c_{n}\partial_{\chi}-\tilde{\mathbb{L}})^{2}$. Notably, this operator is positive semidefinite. An appropriate acceleration operator is thus given by the positive-definite operator ${\bf M}_{n}=r-(\mathbb{B}_{n})^{2}$, where $r>0$ is a positive number whose choice is arrived at via numerical experiments.